Gravitational Lensing in the Kerr Spacetime Geometry Short list of problems/issues

(1) No analytic description of relativistic caustics (except for the intersections on the equatorial plane.

(2) No analytic description of gravitational lensing out of the equatorial plane.

(3) No idea (analytical or numerical) for what happens to higher order caustics.

(4) Is strong field gravitational lensing observable?

(5) Analytical shape of the caustics in Strong Deflection Limit?

(6) Gravitational lensing far from the equatorial plane in the Strong Deflection Limit?

(7) Description of additional images in the Strong Deflection Limit?

(8) What coordinates are best suited for Lensing in Kerr spacetime?

(9) Projects for MAXIM?

(10) Non-disk lensing events for MAXIM?

(11) Gas trajectories in the accretion disk — fully GR code or MHD/Kerr?

(12) Are the time delays the same for Kerr and “Shifted Schwarzschild”? To what order?

(13) Are caustics/critical curves in first order strong field limit equivalent to “shifted Schwarzschild”?

(14) Plotting the caustics from conjugate points: Which photons (light rays) form the caustics?

(15) What does the lightcone look like?

(16) Plot higher order caustic sheets numerically.

(17) Plot primary caustic sheet for observer close to Black hole.

(18) How do caustics look for observers outside the equatorial plane?

(19) How are the astroid caustics formed?

(20) Is the birth process of the astroid caustic sheet for Kerr stable?

(21) Clarify GR definitions of twist/shear/convergence in terms of lensing definition.

(22) Stability of caustics with respect to metric perturbations?

(23) Are all stable caustic sheets lightlike?